# Light-induced localized vortices in multicomponent Bose–Einstein condensates

### v1

@@ Abstract
We study continuous interaction of a trapped two-component Bose–Einstein condensate with light fields in a $\Lambda$-type configuration. Using light beams with orbital angular momentum, we theoretically show how to create a stable, pinned vortex configuration, where the rotating component is confined to the region surrounded by the second, non-rotating component. The atoms constituting this vortex can be localized in volumes much smaller than the volume occupied by the second component. The position of the vortex can be robustly changed by moving the laser beams as long as the beam movement speed is below the speed of sound.
@@ I Introduction
A key challenge in the study of BEC vortices is their controlled creation and manipulation. Typically, the vortices are produced by transferring angular momentum to the atoms using optical means: through phase imprinting [6], stirring [7, 8], or using beams carrying orbital angular momentum, such as Laguerre–Gaussian beams [9, 10, 11, 12, 13, 14, 15, 16, 17]. More recently, interest has also increased in coupling light to condensates using the so-called $\Lambda$-type configurations, which create an internal dark state. The properties of this state can be used to realize atom control at subwavelength resolution [18, 19], create optical lattices featuring barriers of subwavelength width [20, 21], and make narrow structures in the BEC [22]. The $\Lambda$-systems have also been used to create vortices in BEC via time-dependent transfer of population between the atomic internal states by applying Raman-type schemes [11, 23, 12, 24].
In this work, we consider a trapped two-component BEC continuously interacting with two light fields in a $\Lambda$-type configuration. One of the two laser fields used in the setup is an LG beam, which can transfer angular momentum to the BEC atoms. We show in the following that the driven system possesses stationary states such that either one or both components of the BEC are in a vortex state. These stationary states belong to the manifold of dark states and are therefore immune to decay via spontaneous emission. Compared to the familiar coreless-vortex profiles (whereby the core of the rotating vortex is filled by the non-rotating component), the stationary states can have ‘inverted’ profiles, with the rotating component being surrounded by the non-rotating one. In this case, the rotating atoms of one component are tightly localized in a region of the order of the healing length of the other component. The degree of localization can be controlled by tuning the relative strength of the two laser beams. Remarkably, when this relative strength exceeds a certain threshold, this kind of state becomes the lowest-energy state of the dark-state manifold. To further demonstrate the stability of these states, we show that the resulting vortices can be robustly moved around the trap by moving the laser beams. As long as the beam movement speed is below the speed of sound, the vortex follows the beams without notably disturbing the non-rotating component.
@@ II.1 Single-particle Hamiltonian of the system
where $(\rho,\varphi)$ are the polar coordinates, $w_{0}$ is the beam waist (the same for both beams), $\Omega_{0}$ is the amplitude of the Rabi frequencies, and $k_{z}$ is the wave number for the paraxial propagation of the light beams in the $z$ direction. The parameter $a$, having dimension of length, controls the subwavelength nature of the setup. The beam characterized by the Rabi frequency $\Omega_{1}$ couples the atomic internal states $|1\rangle$ and $|3\rangle$ and represents a Laguerre–Gauss mode ${\rm LG}_{0}^{\nu}$ with a vorticity (winding number) $\nu>0$. On the other hand, the atomic states $|2\rangle$ and $|3\rangle$ are coupled by a Gaussian beam characterized by the Rabi frequency $\Omega_{2}$. Both beams co-propagate along the $z$-axis [26, 28, 27], oscillating at frequencies $\omega_{1}$ and $\omega_{2}$, respectively.
@@
The temporal dependence of this Hamiltonian has been eliminated by transitioning to the rotating frame and adopting the rotating wave approximation [30]. Two parameters emerge as a result, including the single-photon detuning $-\varDelta=E_{3}-E_{1}-\hbar\omega_{1}$ and the two-photon detuning $\epsilon=E_{2}-E_{1}+\hbar\omega_{2}-\hbar\omega_{1}$; here $E_{i}$ denotes the energy of the $i$th energy level (see Fig. 1). Furthermore, we have introduced the spontaneous decay rate $\Gamma$ for the excited state $|3\rangle$ represented through the imaginary part of the excited-state energy. In the following, we consider the two-photon resonance, for which $\epsilon=0$.
@@
|±$|\pm\rangle=\frac{1}{\sqrt{2}}\left(|{\rm B}\rangle\pm|3\rangle\right)\,,\quad\mathrm{where}\quad|{\rm B}\rangle=\frac{\zeta^{*}|1\rangle+|2\rangle}{\sqrt{1+|\zeta|^{2}}}\,,$ (6)
@@ II.2 Dynamics of the BEC
Here $|i\rangle$ are the bare atomic states featured in the definition of $\hat{H}_{0}$ in Eq. (3). To account for the interaction between the atoms, we use the Gross–Pitaevskii approach for the multicomponent Bose–Einstein condensates [31]. In practice, this amounts to supplementing the Schrödinger equations for $\Phi_{i}(\boldsymbol{r})$ with the nonlinear terms, thereby promoting the single-particle wave functions to the wave functions (order parameters) of the components of the condensate [27]. This way we arrive at
@@
For atoms adiabatically following the dark state, the population of the excited state $|3\rangle$ described by the wave function $\Phi_{3}$ is small at all times, so atom collisions are not taken into account in the equation for $\Phi_{3}$. The coefficients $g_{ij}$ describing interaction between the atoms in the corresponding internal states are related to the $s$-wave scattering lengths $a_{ij}$ as $g_{ij}=4\pi a_{ij}/m$. The multicomponent wave function of the BEC is normalized to the total number of atoms, $\sum_{i=1}^{3}\int\mathrm{d}\boldsymbol{r}\,|\Phi_{i}|^{2}=N$. We will also use parameters $\eta_{i}$ to quantify the relative population of each component:
@@
Reduction to the dark-state manifold thus results in the appearance of geometric vector and scalar potentials $\boldsymbol{A}(\boldsymbol{r})$ and $U(\boldsymbol{r})$ emerging due to the position-dependence of the atom–light coupling and given by [28, 27]
@@
$\begin{split}\boldsymbol{A}&=-\frac{\hbar\nu}{a}\frac{\left(\frac{\rho}{a}\right)^{2\nu-1}}{1+\left(\frac{\rho}{a}\right)^{2\nu}}\boldsymbol{e}_{\phi},\\ U&=\frac{1}{m}\left(\frac{\hbar\nu}{a}\right)^{2}\frac{\left(\frac{\rho}{a}\right)^{2\nu-2}}{\left[1+\left(\frac{\rho}{a}\right)^{2\nu}\right]^{2}}.\end{split}$ (15)
@@ II.3 Trap geometry and parameter values
As the geometric vector and scalar potentials $\boldsymbol{A}$ and $U$ of Eq. (15), corresponding to the light beams (1), do not depend on $z$, we can consider the two-dimensional motion (2D) of dark-state atoms in the $xy$ plane for fixed $z$. Specifically, for a cylindrically symmetric setup of interest, in the dark-state GPE (13) we can separate the variables as $\Phi_{{\rm D}}(\rho,\varphi,z,t)=\Phi_{{\rm D}}^{(2)}(\rho,\varphi,t)Z(z)$ and take $Z(z)=1/\sqrt{d}$, thereby using the Thomas–Fermi approximation for the atomic motion in the $z$-direction extended over the distance $d$. The resulting equation for $\Phi_{{\rm D}}^{(2)}(\rho,\varphi,t)$ has the same form as Eq. (13) but the interaction strength changes as $\beta\to\beta/d$. A similar procedure can be performed for the exact system (10), yielding accurate results in the adiabatic regime.
@@
[…]
In all our calculations we used $\Omega_{0}=20.0\,\mathrm{MHz}$ and $\varDelta=10\Omega_{0}/a^{\nu}$, where the parameter $a$ characterizes the Rabi frequency of the vortex light beam in Eq. (1). The beam-waist term $\mathrm{e}^{-\rho^{2}/w_{0}^{2}}$ can be set to unity assuming the waist is much larger than the radial extent of the cloud. Moreover, the term $\mathrm{e}^{-\rho^{2}/w_{0}^{2}}$ plays no role in the adiabatic dynamics of atoms in the dark state, since it drops out in the relative Rabi frequency $\zeta$ given by Eq. (5). However, we did include this term in the exact numerical caluculations by setting $w_{0}=2R$ — while the results were the same, this provided a slight speed-up for the calculations.
@@ III.1.2 Small $a$
Now let us consider the case when the distance $a$ is small compared to the system radius, $a\ll R$. In that case, one has $\rho\gg a$ for most of the area occupied by the condensate, except for a small area close to the center, $\rho\lessapprox a$ in which the geometric scalar and vector potentials are concentrated. For $\rho\gg a$, it follows from Eq. (15) that $\rho A_{\varphi}(\rho)\to-\nu$. The ground-state solution can be found by choosing $l$ such that $l+\nu=0$; the solution will again be given by $J_{0}$. Thus, for $\rho\gg a$ the ground state wave function is given by $\psi^{(-\nu)}(\rho,\varphi)=CJ_{0}(\rho j_{0,1})\mathrm{e}^{-\mathrm{i}\nu\varphi}$. However, close to the origin, where $\rho\lessapprox a$, the radial part of the solution will have a different form. Indeed, $l$ is now fixed to $-\nu$, and we can neglect the term $\rho A_{\varphi}$. The solution in the region $\rho\ll a$ is then given by $J_{\nu}$, having a zero at the origin as a vortex solution should. Meanwhile, in the region $\rho\gg a$, the two lowest excited states will come in a degenerate pair having the radial dependence $J_{1}$ and corresponding to $l+\nu=\pm 1$. Close to the origin ($\rho\ll a$), they will have the radial dependencies $J_{\pm 1-\nu}$. These results are illustrated in Figs. 2(b1)–(b5) for the case where $a=0.01$, $\nu=1$. In Fig. 2(b1), it is apparent that the maximum value of $U$ greatly exceeds that of the chemical potential. However, the presence of the potential does not appreciably influence the resulting wave unctions due to the potential being localized at the origin. Figure 2(b2) displays the curve $\rho A_{\varphi}$, which has a value close to $\nu=-1$ in most of space occupied by the condensate. In agreement with the preceding arguments, the state with the lowest chemical potential is obtained for $l=-1$ [see Fig. 2(b3)]. Away from the origin, the wave function tends to $J_{0}$, while at the origin it attains a zero. Further reducing $a$ increases the steepness of the wave function near the origin and enhances the match with $J_{0}$ (not shown). The degenerate pair of the lowest excited states is shown in Figs. 2(b4)–(b5). These wave functions mostly follow the shape of $J_{1}$ except for the small portion of space near the origin where the condition $\rho\gg a$ no longer holds. There, the state with $l=+1-\nu=0$ has the form of a $J_{0}$ function, while the $l=-1-\nu=-2$ state has the shape of $J_{2}$.
Figure 3: The potentials $U$ and $\rho A_{\varphi}$, as well as the solutions of Eq. (10) for $\nu=1$ and $a=0.5$. (a) Potentials $U$ and $\rho A_{\varphi}$. (b), (c), and (d) show, respectively, three lowest-energy solutions of Eq. (10); wave function normalization is $\sum_{i=1}^{3}\int_{S}\,|\psi_{i}|^{2}\,\mathrm{d}S=1$. The upper panels show the radial cuts of the densities $|\psi_{1}|^{2}$ of the first component, while the lower panels show the same for the second component. The values of the density of the third component are vanishingly small in all cases (calculation yields values no larger than $10^{-7}$) and are therefore not shown. For the same reason, the relative occupation of the second component can be taken to be $\eta_{2}=1-\eta_{1}$. The vorticities $l_{i}$ of the two components are displayed together with the insets showing the two-dimensional phase profiles in the $xy$-plane. The value of the phase is color-coded as follows: dark $\text{dark blue}=-\pi$, $\text{white}=0$, dark $\text{dark red}=\pi$.
@@
Thus, if parameter $a$ is small enough so that the condition $\rho\gg a$ holds in most of space occupied by the condensate, the dark-state solution of the lowest energy will be the one corresponding to $l=-\nu$. In this limit the gradient of the phase $\phi$ of wave function equals $-\boldsymbol{e}_{\varphi}\nu/\rho$, thereby canceling the contribution of vector potential $\boldsymbol{A}$ to the velocity field: $\boldsymbol{v}\propto\nabla\phi-\boldsymbol{A}\to 0$. As a result, $\Phi_{1}$ will be in a vortex state, while $\Phi_{2}$ will be vortex-free.

### v2

@@ Abstract
We study continuous interaction of a trapped two-component Bose–Einstein condensate with light fields in a $\Lambda$-type configuration. Using light beams with orbital angular momentum, we theoretically show how to create a stable, pinned vortex configuration, where the rotating component is confined to the region surrounded by the second, non-rotating component. The atoms constituting this vortex can be localized in volumes much smaller than the volume occupied by the second component. We also show that the vortex position can be changed dynamically by moving the laser beams, provided the beams’ movement speed remains below the speed of sound. This allows us to use the localized vortex to stir the second component, and to determine the superfluid flow’s critical velocity.
@@ I Introduction
A key challenge in the study of BEC vortices is their controlled creation and manipulation. Typically, the vortices are produced by transferring angular momentum to the atoms using optical means: through phase imprinting [6], stirring [7, 8], or using beams carrying orbital angular momentum, such as Laguerre–Gaussian (LG) beams [9, 10, 11, 12, 13, 14, 15, 16, 17]. More recently, interest has also increased in coupling light to condensates using the so-called $\Lambda$-type configurations, which create an internal dark state. The properties of this state can be used to realize atom control at subwavelength resolution [18, 19], create optical lattices featuring barriers of subwavelength width [20, 21], and make narrow structures in the BEC [22]. The $\Lambda$-systems have also been used to create vortices in BEC via time-dependent transfer of population between the atomic internal states by applying Raman-type schemes [11, 23, 12, 24].
In this work, we consider a trapped two-component BEC continuously interacting with two light fields in a $\Lambda$-type configuration. One of the two laser fields used in the setup is an LG beam, which can transfer angular momentum to the BEC atoms. We show in the following that the driven system possesses stationary states such that either one or both components of the BEC are in a vortex state. These stationary states belong to the manifold of dark states and are therefore immune to decay via spontaneous emission. Compared to the familiar coreless-vortex profiles (whereby the core of the rotating vortex is filled by the non-rotating component), the stationary states can also have “inverted” profiles, with the rotating component being surrounded by the non-rotating one. In this case, the rotating atoms of one component are tightly localized in a region of the order of the healing length of the other component. The degree of localization can be controlled by tuning the relative strength of the two laser beams. Remarkably, when this relative strength exceeds a certain threshold, the localized-vortex state becomes the lowest-energy state of the dark-state manifold. To further demonstrate the stability of these states, we show that the resulting vortices can be robustly moved around the trap by moving the laser beams. As long as the beam movement speed is below the speed of sound, the vortex follows the beams without notably disturbing the non-rotating component. To quantify this effect, we move the vortex in circular paths, thereby stirring the other component with a localized impurity. By examining the resulting heating, we determine the critical velocity of the superfluid flow below which the vortex moves without dissipation.
@@ II.1 Single-particle Hamiltonian of the system
where $(\rho,\varphi)$ are the polar coordinates, $w_{0}$ is the beam waist (the same for both beams), $\Omega_{0}$ is the amplitude of the Rabi frequencies, and $k_{z}$ is the wave number for the paraxial propagation of the light beams in the $z$ direction. The parameter $a$, having dimension of length, controls the subwavelength nature of the setup. The beam characterized by the Rabi frequency $\Omega_{1}$ couples the atomic internal states $|1\rangle$ and $|3\rangle$ and represents a Laguerre–Gaussian mode ${\rm LG}_{0}^{\nu}$ with a vorticity (winding number) $\nu>0$. On the other hand, the atomic states $|2\rangle$ and $|3\rangle$ are coupled by a Gaussian beam characterized by the Rabi frequency $\Omega_{2}$. Both beams co-propagate along the $z$-axis [26, 28, 27], oscillating at frequencies $\omega_{1}$ and $\omega_{2}$, respectively.
@@
The temporal dependence of this Hamiltonian has been eliminated by transitioning to the rotating frame and adopting the rotating wave approximation [30]. Two parameters emerge as a result, including the single-photon detuning $\varDelta=E_{3}-E_{1}-\hbar\omega_{1}$ and the two-photon detuning $\epsilon=E_{2}-E_{1}+\hbar\omega_{2}-\hbar\omega_{1}$; here $E_{i}$ denotes the energy of the $i$th energy level (see Fig. 1). Furthermore, we have introduced the spontaneous decay rate $\Gamma$ for the excited state $|3\rangle$ represented through the imaginary part of the excited-state energy. In the following, we consider the two-photon resonance, for which $\epsilon=0$.
@@
$|\pm\rangle=\frac{1}{\sqrt{2}}\left(|{\rm B}\rangle\pm|3\rangle\right)\,,\quad\mathrm{where}\quad|{\rm B}\rangle=\frac{\zeta^{*}|1\rangle+|2\rangle}{\sqrt{1+|\zeta|^{2}}}\,$, (6)
@@ II.2 Dynamics of the BEC
Here $|i\rangle$ are the bare atomic states featured in the definition of $\hat{H}_{0}$ in Eq. (3). To account for the interaction between the atoms, we use the Gross–Pitaevskii approach for the multicomponent BECs [31]. In practice, this amounts to supplementing the Schrödinger equations for $\Phi_{i}(\boldsymbol{r})$ with the nonlinear terms, thereby promoting the single-particle wave functions to the wave functions (order parameters) of the components of the condensate [27]. This way we arrive at
@@
For atoms adiabatically following the dark state, the population of the excited state $|3\rangle$ described by the wave function $\Phi_{3}$ is small at all times, so atom collisions are not taken into account in the equation for $\Phi_{3}$. The coefficients $g_{ij}$ describing interaction between the atoms in the corresponding internal states are related to the $s$-wave scattering lengths $a_{ij}$ as $g_{ij}=4\pi\hbar^{2}a_{ij}/m$. The multicomponent wave function of the BEC is normalized to the total number of atoms, $\sum_{i=1}^{3}\int\mathrm{d}\boldsymbol{r}\,|\Phi_{i}|^{2}=N$. We will also use parameters $\eta_{i}$ to quantify the relative population of each component:
@@
It is worth noting that, although $\Phi_{\rm D}$ is the solution to a nonlinear mean-field equation for a BEC, the term “dark-state wave function” remains appropriate. This state contains no atoms in the internal state $|3\rangle$, and, provided the adiabatic approximation holds, this remains true at all times, since the equation for $\Phi_{\rm D}$ is fully decoupled from those for $\Phi_{\rm B}$ and $\Phi_{3}$. In contrast, the equation for the “bright-state BEC” $\Phi_{\rm B}$ is coupled to $\Phi_{3}$, allowing atoms in $\Phi_{\rm B}$ to transition into $|3\rangle$ and decay.
It is apparent from Eq. (13) that reduction to the dark-state manifold results in the appearance of geometric vector and scalar potentials $\boldsymbol{A}(\boldsymbol{r})$ and $U(\boldsymbol{r})$ emerging due to the position-dependence of the atom–light coupling and given by [28, 27]
@@
$\begin{split}\boldsymbol{A}&=-\frac{\hbar\nu}{a}\frac{\left(\frac{\rho}{a}\right)^{2\nu-1}}{1+\left(\frac{\rho}{a}\right)^{2\nu}}\boldsymbol{e}_{\varphi},\\ U&=\frac{1}{m}\left(\frac{\hbar\nu}{a}\right)^{2}\frac{\left(\frac{\rho}{a}\right)^{2\nu-2}}{\left[1+\left(\frac{\rho}{a}\right)^{2\nu}\right]^{2}}.\end{split}$ (15)
@@
We remark that the dark state potentials arising from coupling BECs with LG beams have previously been considered (see, in particular, Ref. [27]). However, the stationary states emerging under continuous interaction with the laser beams have not been analyzed before.
@@ II.3 Trap geometry and parameter values
As the geometric vector and scalar potentials $\boldsymbol{A}$ and $U$ of Eq. (15), corresponding to the light beams (1), do not depend on $z$, we can consider the two-dimensional (2D) motion of dark-state atoms in the $xy$ plane for fixed $z$. Specifically, for a cylindrically symmetric setup of interest, in the dark-state GPE (13) we can separate the variables as $\Phi_{{\rm D}}(\rho,\varphi,z,t)=\Phi_{{\rm D}}^{(2)}(\rho,\varphi,t)Z(z)$ and take $Z(z)=1/\sqrt{d}$, thereby using the Thomas–Fermi approximation for the atomic motion in the $z$-direction extended over the distance $d$. The resulting equation for $\Phi_{{\rm D}}^{(2)}(\rho,\varphi,t)$ has the same form as Eq. (13) but the interaction strength changes as $\beta\to\beta/d$. A similar procedure can be performed for the exact system (10), yielding accurate results in the adiabatic regime.
@@
The typical values of the Rabi frequencies used in $\Lambda$-scheme experiments range from $\sim 10\,\mathrm{MHz}$ to $\sim 1\,\mathrm{GHz}$. Our calculations show that the dark-state GPE yields very accurate results [compared to the solution of the full system of GPEs (10)] already at $\Omega_{0}=1\,\mathrm{MHz}$, meaning that the lowest states of the dark-state manifold indeed become effectively decoupled from the states of the $|\pm\rangle$ manifolds. Generally, the non-adiabatic coupling may be disregarded provided $U/\hbar\ll\Omega(\boldsymbol{r})$ for all $\boldsymbol{r}$ [27, 28, 20], where $\Omega\equiv\Omega(\boldsymbol{r})$ is the total Rabi frequency given by Eq. (7). For $\nu=1$, the maximum of $U$ is located at the origin, while $\Omega$ attains its minimal value there (we neglect the beam-waist term $\mathrm{e}^{-\rho^{2}/w_{0}^{2}}$ in the present reasoning), equal to $\Omega(\boldsymbol{\boldsymbol{r}}=0)=\Omega_{0}$. However, even for the lowest value of $a$ considered in this work ($a=0.3\,\mathrm{\mu m}$), one finds $U/\hbar=8\,\mathrm{kHz}$, which is orders of magnitude smaller than the typical value of $\Omega_{0}$. Furthermore, although generally including the decay rate of $|3\rangle$ improves the validity of the dark-state description [20, 29, 21], in the regime in question we obtained equally accurate results even for $\Gamma=0$ — again because $\Omega_{0}$ is very large compared to $U$. Finally, we mention the effect of the detuning $\varDelta$. Non-zero positive values of $\varDelta$ shift the manifold of the $|$-$\rangle$ states down in energy, further reducing the coupling with the states of the $|{\rm D}\rangle$ manifold. On the other hand, the states of the $|+\rangle$ manifold are then also shifted to lower energies, and this enhances the coupling with the $|{\rm D}\rangle$-manifold states. Thus, the dark-state GPE is expected to be valid for some intermediate values of $\varDelta$. For example, we have found that for $\Omega_{0}\sim 10\,\mathrm{MHz}$, values of $\varDelta$ from 0 to up to $\sim 10\Omega_{0}/a^{\nu}$ lead to good agreement between the dark-state analysis and the full treatment. Meanwhile, negative values of $\varDelta$ shift to higher energies the manifold of $|$-$\rangle$ states. This facilitates coupling between the $|$-$\rangle$ and $|{\rm D}\rangle$ states, reducing the accuracy of the dark-state description.
In all our calculations we used $\Omega_{0}=20.0\,\mathrm{MHz}$ and $\varDelta=10\Omega_{0}/a^{\nu}$, where the parameter $a$ characterizes the Rabi frequency of the vortex light beam in Eq. (1). The beam-waist term $\mathrm{e}^{-\rho^{2}/w_{0}^{2}}$ can be set to unity assuming the waist is much larger than the radial extent of the cloud. Moreover, the term $\mathrm{e}^{-\rho^{2}/w_{0}^{2}}$ plays no role in the adiabatic dynamics of atoms in the dark state, since it drops out of the relative Rabi frequency $\zeta$ given by Eq. (5). However, we did include this term in the exact numerical calculations by setting $w_{0}=2R$ — while the results were the same, this provided a slight speed-up for the calculations.
@@ III.1.2 Small $a$
Now let us consider the case when the distance $a$ is small compared to the system radius, $a\ll R$. In that case, one has $\rho\gg a$ for most of the area occupied by the condensate, except for a small area close to the center, $\rho\lessapprox a$ in which the geometric scalar and vector potentials are concentrated. For $\rho\gg a$, it follows from Eq. (15) that $\rho A_{\varphi}(\rho)\to-\nu$. The ground-state solution can be found by choosing $l$ such that $l+\nu=0$; the solution will again be given by $J_{0}$. Thus, for $\rho\gg a$ the ground state wave function is given by $\psi^{(-\nu)}(\rho,\varphi)=CJ_{0}(\rho j_{0,1})\mathrm{e}^{-\mathrm{i}\nu\varphi}$. However, close to the origin, where $\rho\lessapprox a$, the radial part of the solution will have a different form. Indeed, $l$ is now fixed to $\nu$, and we can neglect the term $\rho A_{\varphi}$. The solution in the region $\rho\ll a$ is then given by $J_{\nu}$, having a zero at the origin as a vortex solution should. Meanwhile, in the region $\rho\gg a$, the two lowest excited states will come in a degenerate pair having the radial dependence $J_{1}$ and corresponding to $l+\nu=\pm 1$. Close to the origin ($\rho\ll a$), they will have the radial dependencies $J_{\pm 1-\nu}$. These results are illustrated in Figs. 2(b1)–(b5) for the case where $a=0.01$, $\nu=1$. In Fig. 2(b1), it is apparent that the maximum value of $U$ greatly exceeds that of the chemical potential. However, the presence of the potential does not appreciably influence the resulting wave functions due to the potential being localized at the origin. Figure 2(b2) displays the curve $\rho A_{\varphi}$, which has a value close to $\nu=-1$ in most of space occupied by the condensate. In agreement with the preceding arguments, the state with the lowest chemical potential is obtained for l=1l=-1 [see Fig. 2(b3)]. Away from the origin, the wave function tends to $J_{0}$, while at the origin it attains a zero. Further reducing $a$ increases the steepness of the wave function near the origin and enhances the match with $J_{0}$ (not shown). The degenerate pair of the lowest excited states is shown in Figs. 2(b4)–(b5). These wave functions mostly follow the shape of $J_{1}$ except for the small portion of space near the origin where the condition $\rho\gg a$ no longer holds. There, the state with $l=+1-\nu=0$ has the form of a $J_{0}$ function, while the $l=-1-\nu=-2$ state has the shape of $J_{2}$.
@@
Thus, if parameter $a$ is small enough so that the condition $\rho\gg a$ holds in most of space occupied by the condensate, the dark-state solution of the lowest energy will be the one corresponding to $l=-\nu$. In this limit the gradient of the phase $\phi$ of wave function equals $\boldsymbol{e}_{\varphi}\nu/\rho$, thereby canceling the contribution of vector potential $\boldsymbol{A}$ to the velocity field: $\boldsymbol{v}\propto\nabla\phi-\boldsymbol{A}\to 0$. As a result, $\Phi_{1}$ will be in a vortex state, while $\Phi_{2}$ will be vortex-free.